# Selective Variable Convolution Meets Dynamic Content Guided Attention for Infrared Small Target Detection

### v1

@@ III-C1 Stage I: Coarse-Grained Attention with Joint Channel-Space Fusion
where $W_{\text{coa}}\in\mathbb{R}^{C\times H\times W}$ is rough SIM, $\oplus$ denotes element-wise addition with automatic dimension expansion.
Figure 4: The proposed structure of the DCGA: employs a parallel approach to equally compute the weights of spatial and channel information. It also incorporates random mixing of channel and spatial information to enhance the capability for target segmentation.
III-C2 Stage II: Fine-Grained Calibration with Dynamic SIM Generation
The fine-grained calibration include channel shuffling fusion and dynamic weight generation.
Channel Shuffling Fusion: The raw feature $X\in\mathbb{R}^{C\times H\times W}$ and coarse attention map $W_{coa}$ are concatenated and then shuffled along the channel dimension [46], establishing inter-channel correlations while preserving spatial details.
Dynamic Weight Generation: A gated convolution layer $\mathcal{GC}_{7\times 7}$ followed by sigmoid activation generates channel-specific SIM :
$W=\sigma\left(\operatorname{\mathcal{GC}}_{7\times 7}\left(\operatorname{CS}\left(\operatorname{Concat}\left(X,W_{\text{coa}}\right)\right)\right)\right)$ (8)
where $\operatorname{CS}(\cdot)$ denotes the channel shuffling operation. This enables content-adaptive calibration by allowing each channel’s SIM to dynamically adjust based on local feature patterns.
DCGA assigns a unique SIM to each channel, guiding the model to focus on the important areas of each channel. Therefore, more useful information encoded in the features can be emphasized to effectively improve IRSTD.
III-D Adaptive Dynamic Feature Fusion Module
The proposed Adaptive Dynamic Feature Fusion(ADFF) module aims to enhance the network’s ability to distinguish between actual targets and similar false alarms. Specifically, we introduced a dynamic space selection mechanism within the original skip connections, as shown in Fig. 5. This mechanism leverages multi-scale features to enable the network to focus on the areas with the richest spatial contextual information, thereby increasing sensitivity to the target area. The ADFF module is specifically designed to process the feature maps of the encoder-decoder pair in stage i, where the feature maps are represented as:
$X_{i},X^{\prime}_{i}\in\mathbb{R}^{C_{i}\times\frac{H}{2^{i}}\times\frac{W}{2^{i}}}$ (9)
where $C_{i}$ denotes channel number, $\frac{H}{2^{i}}$ and $\frac{W}{2^{i}}$ are height and width of feature maps.
Then we applied channel-wise average pooling and max pooling operations to feature maps $X_{i}$and $X^{\prime}_{i}$, represented as $P_{avg}(\cdot)$and $P_{max}(\cdot)$, respectively.
@@ III-D Adaptive Dynamic Feature Fusion Module
$AD_{i}^{\text{max}}$ $=\operatorname{MaxPool}\left(\operatorname{Concat}\left[X_{i},X^{\prime}_{i}\right]\right).$ (11)
where $AvgPool$ is average pooling operation, $MaxPool$ is max pooling operation.
These two different pooling operations extract the average and maximum feature descriptors of the space, and the generated spatial pooling features are aggregated into a matrix to facilitate information exchange between different spatial descriptors. Next, we use the convolutional layer $F_{2\to N}(\cdot)$to transform these aggregated features into $N$spatial attention maps. These spatial attention maps can effectively guide the subsequent feature fusion process, highlight salient regions, and suppress redundant background information.
Figure 5: The proposed structure of ADFF: integrates the encoding and decoding stages of the network through a spatially adaptive selection mechanism. It utilizes both maximum and average pooling across channels to achieve precise target extraction.
@@
where $\sigma$ denotes the sigmoid function, $\operatorname{Concat}(\cdot)$ is the channel-wise concatenation, and $\operatorname{Conv}_{k\times k}(\cdot)$ represents the convolutional layer (e.g., with a $7\times 7$ kernel, typically outputting a single channel) used to generate the attention map.
The final output $Y_{i}$ of the ADFF module is obtained by first applying the generated attention map $AD_{\text{fuse}}$ to weight the input features $X_{i}$ and $X^{\prime}_{i}$ via element-wise multiplication, and then fusing the results through element-wise addition. This process is formulated as:
$Y_{i}=X_{i}\odot AD_{\text{fuse}}+X^{\prime}_{i}\odot AD_{\text{fuse}}$ (13)
where $\odot$ denotes element-wise multiplication, $Y_{i}$ is the output of ADFF.
The fusion process generates dynamic fusion weights, which can adaptively adjust the importance of each feature map in the final decision. Through this adaptive selective fusion strategy, the ADFF module can automatically identify and integrate the most discriminative features, thereby improving the network’s representation ability of IRSTD.
IV Experiment
@@
table
TABLE I: Configurations for all comparative experiments
Methods(Source & Year) Key parameters configurations
Background Suppression Methods
NWMTH [12]PR’2010 $R_{o}=9,R_{i}=4$
FKRW [47]TGRS’2019 Windows size: 11, $K=4,p=6,\beta=200$
HVS-based Methods
MPCM [17]PR’2016 $N=1,3,\cdots,9$, threshold $k=3$
WLDM [48]TGRS’2016 $L=4,m=2,n=2$, threshold $k=2$
RLCM [49]GRSL’2018 $k_{1}=[2,5,9],k_{2}=[4,9,16]$, scale: 3, threshold $k=1$
ILCM [50]GRSL’2020 Cell size: $3\times 3$, threshold $k=3$
TLLCM [41]GRSL’2020 Scales:[5,7,9] $\lambda=0.5$, threshold $k=3$
GSWLCM [18]GRSL’2022 Local Window Structures: $[3,5,7,9]$, $\delta=0.01,k=20$
Low-rank and Sparse Decomposition
IPI [42]TIP’2013 Patch size: 50, sliding step: 10, $\lambda=1/\sqrt{\mathrm{min}(m,n)}$, $\epsilon=10^{-7}$
RIPT [24]JSTARS’2017 Patch size: 50, sliding step: 10, $\lambda=2/\sqrt{\mathrm{min}(m,n)}$, $\epsilon=10^{-2}$, $\omega=10^{-7}$
PSTNN [26]RS’2019 Patch size: 40, sliding step: 40, $\lambda=0.6/\sqrt{\mathrm{max}(n_{1},n_{2})\times n_{3}}$, $\epsilon=10^{-7}$
Deep Learning Methods
ACM [32]WACV’2021 Backbone: FPN, layer blocks: [4, 4, 4], channels: [8, 16, 32, 64]
AGPCNet [38]TAES’2023 Backbone: resnet34, scales: [10, 6, 5, 3], reduce ratios: [16, 4], gca type: patch, gca att: post
DNANet [34]TIP’2022 Backbone: resnet18, layer blocks: [2, 2, 2, 2], filter: [16, 32, 64, 128, 256]
ISNet [31]CVPR’2022 Backbone: resnet18, layer blocks: [4, 4, 4], channels: [8, 16, 32, 64], fuse mode: AsymBi
UIUNet [33]TIP’2023 Channels: [64, 128, 256, 512], fuse mode: AsymBi
MTUNet [51]TGRS’2023 Backbone: resnet18, layer blocks: [2, 2, 2, 2], filter: [16, 32, 64, 128, 256]
SeRankDet TGRS’2024 Channels:[64, 128, 256, 512, 1024],o=3
SCTransNet [52] TGRS’2024 Backbone:Swin-Tiny, Layer Num=[2,6,6,2], Head Num=[3,6,12,24]
@@ IV-A1 Datasets
To rigorously evaluate the performance and generalization capabilities of our method, we conduct comprehensive benchmark experiments on three widely recognized infrared small target detection datasets: NUAA-SIRST [32], IRSTD-1K [31], and NUDT-SIRST [34]. These datasets demonstrate considerable diversity in target characteristics, background clutter, and imaging conditions, thereby enabling robust validation of detection methodologies within real-world operational constraints.
@@
    NUAA-SIRST: The real long-range infrared single-frame dataset comprises 427 images capturing diverse realistic scenes, such as clouds, cities, and oceans. However, the dataset is relatively small in scale, with only 480 target instances, which may limit its effectiveness for training and evaluating models in scenarios that demand high data diversity.
@@
    IRSTD-1K: IRSTD-1K is designed to address the challenges of multi-scene cross-domain detection, comprising 1,001 highly diverse images. The complexity of its scene types and target shapes imposes significantly higher demands on the model’s generalization capabilities compared with the other two datasets.
@@
    NUDT-SIRST: The scene-enhanced dataset, constructed from synthetic data, comprises 1,327 images with complex backgrounds, including cities, fields, bright regions, oceans, and clouds. By employing adaptive target size and intensity blurring functions, it generates diverse targets in the form of point-like, patchy, and extended shapes, yielding distribution characteristics that closely resemble those of real-world datasets.
IV-A2 Evaluation Metrics
To quantitative analyze our DCGANet and other methods, we adopt four standard metrics to evaluate detection performance: Intersection over Union (IoU), normalized IoU (nIoU), Probability of detection (Pd), and False alarm rate (Fa). The detailed description of these metrics are as follows:
  1. 1.
    IoU measures the accuracy of object localization by comparing the overlap between the predicted bounding box and the ground truth. It is defined as the ratio of the intersection area to the union area of the two bounding boxes.
    $\text{IoU}=\frac{TP}{T+P-TP}$ (14)
  2. 2.
    nIoU is a standardized adaptation of the IoU metric.
    $\text{nIoU}=\frac{1}{N}\sum_{i=1}^{N}\frac{TP(i)}{T(i)+P(i)-TP(i)}$ (15)
  3. 3.
    Pd quantifies the likelihood of correctly identifying true targets.
    $\text{Pd}=\frac{TP}{TP+FN}$ (16)
  4. 4.
    Fa measures the frequency of incorrect target detections.
    $\text{Fa}=\frac{FP}{FP+TN}$ (17)
where $N$ is the total number of samples, $T$ and $P$ correspond to the number of ground truth and predicted positive pixels, respectively. Additionally, $TP$, $FP$, and $FN$ denote the counts of true positive, false positive, and false negative pixels, respectively.
In addition, the ROC curve illustrates the variation trend of the Probability of detection (Pd) across different False alarm rate (Fa).
IV-A3 Implementation Details
We adopt U-Net as the baseline model, Soft-IoU Loss [53] as the loss function, and AdamW [54] as the optimizer, and incorporating a polynomial learning rate decay strategy to optimize the training process. Additionally, we employ a deep supervision strategy to further enhance model performance. All experiments are conducted using distributed training on a system equipped with four RTX 3090 GPUs. To ensure comprehensive and reliable evaluation, we train independent models on each of the three datasets. The detailed hyperparameter configurations for all comparative methods are summarized in Tab. I. To address the significant differences in dataset distribution and resolution, we customize the hyperparameters for each dataset, with specific configurations provided in Tab. II.
TABLE II: Customized Hyperparameters for Each Dataset
Dataset Epochs Lr Batch Resolution
NUAA-SIRST 500 0.0001 4 512$\times$512
IRSTD-1K 500 0.0001 4 512$\times$512
NUDT-SIRST 500 0.0001 16 256$\times$256
TABLE III: Comparison with Other State-of-the-art methods on four datasets. The $\uparrow$ indicates that the higher the indicator, the better, and the $\downarrow$ indicates that the lower the indicator, the better. The best performance is highlighted in bold, and the second-best is underlined.
Method NUAA-SIRST(Tr=60%) IRSTD-1K(Tr=60%) NUDT-SIRST(Tr=60%)
IoU $\uparrow$ nIoU $\uparrow$ $P_{d}$uparrow $F_{a}$downarrow IoU $\uparrow$ nIoU $\uparrow$ $P_{d}$uparrow $F_{a}$downarrow IoU $\uparrow$ nIoU $\uparrow$ $P_{d}$uparrow $F_{a}$downarrow
                      Background Suppression Methods
NWMTH [12] 15.35 17.00 71.10 55.21 18.31 16.74 49.97 21.06 18.87 16.22 62.84 32.45
FKRW [47] 22.10 28.81 81.41 16.67 10.12 16.83 69.88 24.19 10.04 16.76 72.31 66.90
HVS-based Methods
MPCM [17] 25.64 27.63 68.01 45.07 20.72 23.26 58.78 31.10 27.40 38.95 55.54 11.82
WLDM [48] 5.55 3.37 64.97 21.22 8.16 10.32 54.85 14.79 8.19 11.63 16.30 64.14
RLCM [49] 25.56 27.46 86.34 62.16 16.46 21.83 70.86 67.70 8.21 11.42 70.22 31.67
ILCM [50] 18.42 19.16 38.00 40.63 13.67 11.34 43.32 47.67 5.16 6.48 26.27 88.16
TLLCM [41] 18.20 27.62 82.23 16.18 10.86 18.01 68.95 24.99 13.44 16.98 77.02 69.66
GSWLCM [18] 15.67 13.82 72.48 21.96 12.81 13.46 70.57 13.82 21.08 19.92 69.06 42.78
Low-rank and Sparse Decomposition
IPI [42] 26.67 32.57 83.55 11.97 27.92 30.12 81.27 26.18 25.16 34.64 76.26 43.41
RIPT [24] 11.85 14.91 79.48 22.67 14.71 17.43 79.55 28.35 24.13 33.98 78.54 56.24
PSTNN [26] 22.72 23.50 77.36 29.99 28.83 29.09 82.68 39.48 21.26 24.22 72.78 61.26
Deep Learning Methods
ACM [32] 73.38 74.66 96.06 9.94 60.38 59.83 91.82 21.49 73.86 72.97 97.91 76.22
AGPCNet [38] 77.86 75.23 98.92 34.95 68.30 66.18 94.29 15.55 88.55 88.08 97.40 34.09
DNANet [34] 78.82 79.72 98.63 13.37 69.96 69.35 92.02 23.19 91.32 90.10 97.55 30.71
ISNet[31] 80.83 78.55 98.78 4.93 68.74 61.98 95.80 14.97 88.02 88.76 93.14 28.24
UIUNet [33] 79.76 78.74 98.69 4.41 65.20 64.57 92.33 13.45 88.86 88.11 98.40 23.17
MTUNet [51] 78.41 79.09 98.02 8.66 67.04 68.24 93.18 12.86 87.39 87.34 98.15 29.43
SCTransNet [52] 80.88 79.45 96.48 6.48 71.22 70.11 94.45 13.94 92.39 92.32 97.71 28.76
SeRankDet [55] 81.26 79.66 98.74 2.77 74.45 70.03 95.12 10.65 94.16 93.34 98.71 22.36
$\star$ DCGANet (Ours) 82.03 79.81 98.88 2.99 75.59 72.16 96.64 9.21 94.98 94.46 98.75 20.34
IV-B Comparison With State-of-the-Arts
Our DCGANet method is rigorously compared against state-of-the-arts methods across three datasets using metrics include IoU, nIoU, Pd, and Fa. The BS-based methods NWMTH [12], FKRW [47]. The HVS-based Methods MPCM [17], WLDM [48], RLCM [49], ILCM [50], TLLCM [41], GSWLCM [18]. The LRM-based methods IPI [42], RIPT [24], PSTNN [26]. The deep learning methods ACM [32], AGPCNet [38], DNANet [34], ISNet [31], UIUNet [33], MTUNet [51], SCTransNet [52], SeRankDet [55]. The experimental results summarized in Tab. III, including three critical observations:
1) Our proposed DCGANet exhibits superior performance compared to state-of-the-art methods across multiple datasets. Specifically designed to tackle the challenges posed by low signal-to-noise ratio (SNR) and blurred target boundaries in infrared small target detection, our approach incorporates a multi-modal feature extractor that precisely extracts target features while effectively suppressing background clutter. Additionally, a dedicated feature fusion module is included to aggregate target details from various hierarchical levels. This innovative design substantially enhances detection accuracy and robustness, resulting in significant advancements in the field of infrared small target detection.
2) The experimental results reveal a performance gap exceeding 15% in Intersection over Union (IoU) between Deep Learning-based methods and traditional approaches, highlighting the inherent limitations of model-driven techniques that rely on global sparsity or contrast priors. In contrast, data-driven DL methods demonstrate superior generalization capabilities by leveraging semantic feature extraction and deeper hierarchical representations. This allows them to significantly reduce sensitivity to hyperparameters in unstructured scenarios, resulting in more robust and adaptable performance.
3) On the NUAA-SIRST dataset, DCGANet surpasses SeRankDet across all evaluation metrics except for the False alarm rate (Fa). While SeRankDet shows moderate effectiveness in combining global and local features, its performance significantly deteriorates in cluttered environments. In contrast, DCGANet addresses these limitations by leveraging precise target localization and dynamic feature fusion, resulting in superior pixel-level segmentation accuracy.
@@ IV-B Comparison With State-of-the-Arts
Furthermore, Fig. 6 illustrates receiver operating characteristic (ROC) curves for three datasets. The analysis corroborates the numerical results, confirming DCGANet’s stability and superiority over conventional techniques.
@@
Figure 6: ROC curves of different methods on three datasets. Our DCGANet achieves the highest Pd at very low Fa.

### v2

@@ III-B1 Stage 1: Coarse-Grained Attention and Joint Channel-Spatial Fusion
where $W_{\text{coa}}\in\mathbb{R}^{C\times H\times W}$, and $\oplus$ denotes element-wise addition with automatic dimension expansion.
III-B2 Stage 2: Fine-Grained Calibration for Dynamic SIM Generation
This stage refines the coarse attention map.
Channel-Shuffle Fusion: The original features $X\in\mathbb{R}^{C\times H\times W}$ and the coarse attention map $W_{coa}$ are concatenated and then shuffled along the channel dimension[35], thereby establishing cross-channel correlations while preserving spatial details.
Dynamic Weight Generation: A gated convolutional layer $\mathcal{GC}_{7\times 7}$ followed by a Sigmoid activation function generates the final channel-specific SIM $W$:
$W=\sigma\left(\operatorname{\mathcal{GC}}_{7\times 7}\left(\operatorname{CS}\left(\operatorname{Concat}\left(X,W_{\text{coa}}\right)\right)\right)\right),$ (8)
where $\operatorname{CS}(\cdot)$ denotes the channel shuffle operation and $\sigma(\cdot)$ is the Sigmoid function. This content-adaptive calibration allows the SIM for each channel to be dynamically adjusted based on local feature patterns, guiding the model to focus on important regions within each channel.
Figure 5: The structure of the ADFF module: It integrates the encoding and decoding stages of the network via a spatially adaptive selection mechanism. It utilizes cross-channel max-pooling and average-pooling for precise target extraction.
III-C Adaptive Dynamic Feature Fusion
Our proposed Adaptive Dynamic Feature Fusion (ADFF) module is designed to enhance the network’s ability to distinguish real targets from similar false positives. As shown in Figure 5, we introduce a dynamic spatial selection mechanism into the original skip connection, leveraging multi-scale features to enable the network to focus on regions richest in spatial context. The ADFF module is specifically designed to process the encoder features $X_{i}$ and decoder features $X^{\prime}_{i}$ at stage $i$:
$X_{i},X^{\prime}_{i}\in\mathbb{R}^{C_{i}\times\frac{H}{2^{i}}\times\frac{W}{2^{i}}},$ (9)
where $C_{i}$ is the number of channels, and $\frac{H}{2^{i}}$ and $\frac{W}{2^{i}}$ are the height and width of the feature maps.
We first apply channel-wise average-pooling and max-pooling operations to the concatenated feature map:
@@ III-C Adaptive Dynamic Feature Fusion
$AD_{i}^{\text{max}}$ $=\operatorname{MaxPool}\left(\operatorname{Concat}\left[X_{i},X^{\prime}_{i}\right]\right),$ (11)
where $\operatorname{AvgPool}$ and $\operatorname{MaxPool}$ are the average-pooling and max-pooling operations, respectively.
Next, we concatenate these two pooled features and pass them through a convolutional layer and a Sigmoid function to generate the final spatial attention map $AD_{\text{fuse}}$:
@@
where $\operatorname{Conv}_{k\times k}$ is a convolutional layer with a kernel size of $k\times k$ (e.g., $7\times 7$).
The final output $Y_{i}$ of the ADFF module is obtained by element-wise multiplication of the generated attention map $AD_{\text{fuse}}$ with the input features $X_{i}$ and $X^{\prime}_{i}$ respectively, and then adding the results:
$Y_{i}=(X_{i}\odot AD_{\text{fuse}})+(X^{\prime}_{i}\odot AD_{\text{fuse}}),$ (13)
where $\odot$ denotes element-wise multiplication. Through this adaptive selective fusion strategy, the ADFF module can automatically identify and integrate the most discriminative features, thereby enhancing the network’s representation capability in infrared small target detection.
TABLE I: Configurations for all comparative experiments
@@
[…]
FKRW [37] (TGRS’2019) Windows size: 11, $K=4,p=6,\beta=200$
[…]
WLDM [38](TGRS’2016) $L=4,m=2,n=2$, threshold $k=2$
RLCM [39](GRSL’2018) $k_{1}=[2,5,9],k_{2}=[4,9,16]$, scale: 3, threshold $k=1$
ILCM [40](GRSL’2020) Cell size: $3\times 3$, threshold $k=3$
TLLCM [27](GRSL’2020) Scales:[5,7,9] $\lambda=0.5$, threshold $k=3$
GSWLCM [12](GRSL’2022) Local Window Structures: $[3,5,7,9]$, $\delta=0.01,k=20$
[…]
RIPT [13] (JSTARS’2017) Patch size: 50, sliding step: 10, $\lambda=2/\sqrt{\mathrm{min}(m,n)}$, $\epsilon=10^{-2}$, $\omega=10^{-7}$
PSTNN [28](RS’2019) Patch size: 40, sliding step: 40, $\lambda=0.6/\sqrt{\mathrm{max}(n_{1},n_{2})\times n_{3}}$, $\epsilon=10^{-7}$
[…]
AGPCNet [22](TAES’2023) Backbone: resnet34, scales: [10, 6, 5, 3], reduce ratios: [16, 4], gca type: patch, gca att: post
DNANet [19](TIP’2022) Backbone: resnet18, layer blocks: [2, 2, 2, 2], filter: [16, 32, 64, 128, 256]
ISNet [32](CVPR’2022) Backbone: resnet18, layer blocks: [4, 4, 4], channels: [8, 16, 32, 64], fuse mode: AsymBi
UIUNet [18](TIP’2023) Channels: [64, 128, 256, 512], fuse mode: AsymBi
MTUNet [41](TGRS’2023) Backbone: resnet18, layer blocks: [2, 2, 2, 2], filter: [16, 32, 64, 128, 256]
SeRankDet [42](TGRS’2024) Channels:[64, 128, 256, 512, 1024],o=3
SCTransNet [43](TGRS’2024) Backbone:Swin-Tiny, Layer Num=[2,6,6,2], Head Num=[3,6,12,24]
@@ IV-A1 Datasets
To rigorously evaluate the performance and generalization capabilities of our method, we conduct comprehensive benchmark experiments on three widely recognized infrared small target detection datasets: NUAA-SIRST, IRSTD-1K, and NUDT-SIRST. These datasets exhibit considerable diversity in target characteristics, background clutter, and imaging conditions, thereby enabling robust validation of detection methodologies.
  • NUAA-SIRST: This dataset contains 427 real-world images capturing diverse scenes, such as clouds, cities, and oceans. With 480 target instances, it serves as a foundational benchmark for evaluating detection performance in common scenarios.
  • IRSTD-1K: Comprising 1,001 highly diverse images, IRSTD-1K is designed to challenge a model’s cross-domain generalization. The complexity of its scene types and target shapes imposes significantly higher demands compared to other datasets.
  • NUDT-SIRST: This synthetic dataset includes 1,327 images with complex backgrounds. By employing adaptive functions for target size and intensity blurring, it generates a wide variety of targets (point-like, patchy, extended), creating a distribution that closely mimics real-world data and tests model robustness against complex interference.
IV-A2 Evaluation Metrics
To quantitatively analyze our DCGANet and the comparative methods, we adopt four standard metrics: Intersection over Union (IoU), normalized IoU (nIoU), Probability of Detection ($P_{d}$), and False-alarm Rate ($F_{a}$).
  • IoU measures the localization accuracy by comparing the overlap between the predicted bounding box and the ground truth. It is defined as the ratio of their intersection to their union.
    $\text{IoU}=\frac{TP}{T+P-TP}$ (14)
  • nIoU is a normalized adaptation of the IoU metric, averaging the IoU over all target-containing images to provide a dataset-level performance score.
    $\text{nIoU}=\frac{1}{N}\sum_{i=1}^{N}\frac{TP(i)}{T(i)+P(i)-TP(i)}$ (15)
  • $P_{d}$ (also known as Recall) quantifies the model’s ability to correctly identify true targets. A higher value is better.
    $P_{d}=\frac{TP}{TP+FN}$ (16)
  • $F_{a}$ measures the rate of pixels incorrectly identified as targets. It is calculated per image, and a lower value is better.
    $F_{a}=\frac{FP}{\text{Number of Pixels in Image}}$ (17)
where $N$ is the total number of images with targets, $T$ and $P$ are the sets of ground-truth and predicted positive pixels, respectively. $TP$, $FP$, and $FN$ denote the counts of true positive, false positive, and false negative pixels. In addition, the Receiver Operating Characteristic (ROC) curve illustrates the trade-off between $P_{d}$ and $F_{a}$.
IV-A3 Implementation Details
We adopt U-Net as the baseline, Soft-IoU Loss as the loss function, and AdamW as the optimizer, incorporating a polynomial learning rate decay strategy. A deep supervision strategy is also employed to enhance performance. All experiments are conducted using distributed training on a system with four NVIDIA RTX 3090 GPUs. To ensure a fair and comprehensive evaluation, we train independent models for each of the three datasets. The hyperparameter configurations for all comparative methods are summarized in Table III-C. Due to significant differences in data distribution and resolution, we customize the training hyperparameters for each dataset, as detailed in Table II.
TABLE II: Customized Training Hyperparameters for Each Dataset.
[…]
IRSTD-1K 500 1e-4 4 $512\times 512$
NUDT-SIRST 500 1e-4 16 $256\times 256$
[…]
IV-B Comparison with State-of-the-Art Methods
We rigorously benchmark our DCGANet against a wide range of state-of-the-art (SOTA) methods, including traditional model-driven techniques (Background Suppression, HVS-based, and Low-rank Matrix-based) and various advanced deep learning models. The comprehensive results on all three datasets are summarized in Table IV-B, leading to several key observations:
@@ IV-B Comparison with State-of-the-Art Methods
    Superior Overall Performance: Our proposed DCGANet consistently achieves the best or second-best performance across all metrics on all three datasets. This demonstrates its superior accuracy and robustness in handling the challenges of low signal-to-noise ratio (SNR) and indistinct target boundaries. The combination of its dynamic feature extraction and fusion modules proves highly effective.
@@
    Data-Driven vs. Model-Driven: A significant performance gap (often >15% in IoU) exists between deep learning (DL) methods and traditional approaches. This highlights the inherent limitations of model-driven techniques, which rely on fixed priors like local contrast or global sparsity. In contrast, data-driven DL methods show superior generalization by learning deep hierarchical representations, making them more adaptable and less sensitive to hyperparameters in complex, unstructured scenes.
@@
    Advantage over SOTA DL Methods: On the challenging NUAA-SIRST and IRSTD-1K datasets, DCGANet surpasses recent powerful models like SeRankDet. For instance, on NUAA-SIRST, DCGANet achieves the highest IoU, nIoU, and $P_{d}$, with only a marginal trade-off in the $F_{a}$ rate compared to SeRankDet. This indicates that while other models may be effective, DCGANet’s architecture provides a better balance of high detection probability and precise pixel-level localization, especially in cluttered environments.
@@
Furthermore, Figure 6 illustrates the Receiver Operating Characteristic (ROC) curves for the three datasets. The analysis visually corroborates the quantitative results in Table IV-B, confirming that DCGANet consistently achieves a higher probability of detection ($P_{d}$) at any given false-alarm rate ($F_{a}$), underscoring its stability and superiority.
TABLE III: Quantitative comparison with state-of-the-art methods on three datasets. For all metrics except $F_{a}$, higher is better ($\uparrow$). For $F_{a}$, lower is better ($\downarrow$). The best performance is highlighted in red, and the second-best is in blue.
[…]
FKRW [37] 21.23 27.67 78.18 16.66 9.72 16.16 67.11 24.18 9.64 16.09 69.44 66.88
[…]
WLDM [38] 5.33 3.23 62.40 21.21 7.84 9.91 52.68 14.79 7.87 11.17 15.65 64.11
RLCM [39] 24.55 26.37 82.92 62.13 15.81 20.96 68.05 67.67 7.89 10.97 67.44 31.65
ILCM [40] 17.69 18.40 36.50 40.61 13.13 10.89 41.60 47.65 4.96 6.22 25.23 88.20
TLLCM [27] 17.48 26.53 78.98 16.17 10.43 17.30 66.22 24.98 12.91 16.31 73.97 69.63
GSWLCM [12] 15.05 13.27 69.61 21.95 12.30 12.93 67.78 13.81 20.25 19.13 66.33 42.77
[…]
RIPT [13] 11.38 14.32 76.33 22.66 14.13 16.74 76.40 28.34 23.18 32.63 75.43 56.21
PSTNN [28] 21.82 22.57 74.29 30.00 27.69 27.94 79.41 39.46 20.41 23.26 69.89 61.24
[…]
AGPCNet [22] 74.77 72.26 95.00 34.94 65.59 63.56 90.55 15.54 85.04 84.59 93.54 34.07
DNANet [19] 75.70 76.57 94.73 13.37 67.19 66.60 88.38 23.18 87.70 86.53 93.69 30.69
ISNet [32] 77.63 75.44 94.86 4.93 66.02 59.53 92.00 14.96 84.53 85.26 89.45 28.22
UIUNet [18] 76.60 75.62 94.79 4.41 62.61 62.01 88.67 13.45 85.34 84.62 94.50 23.16
MTUNet [41] 75.30 75.96 94.14 8.65 64.39 65.54 89.49 12.86 83.93 83.89 94.27 29.42
SCTransNet [43] 77.67 76.30 92.66 6.48 68.39 67.34 90.71 13.94 88.73 88.66 93.84 28.75
SeRankDet [42] 78.04 76.51 94.82 2.77 71.50 67.26 91.36 10.64 90.43 89.64 94.81 22.35
[…]
@@
Figure 6: ROC curves of different methods on the three datasets. Our DCGANet consistently achieves the highest $P_{d}$ at very low $F_{a}$ levels, demonstrating its superior trade-off between detection rate and false alarms.